# Color-avoiding Ramsey for directed paths and vector sequences

### v1

@@ 1.1. Sequences of vectors
Our first result answers this question in a very strong form: when r=q1r=q-1 and $q$ is large, we show that the behavior of both $F_{q,q-1}(n)$ and $G_{q,q-1}(n)$ are close to the lower bound in (4): the true exponent tends to $1$ as $q$ grows.
Theorem 1.2.
For $q\geqslant 4$ and any $n$, we have
$F_{q,q-1}(n)\leqslant G_{q,q-1}(n)\leqslant n^{1+O(1/\sqrt{\log q})}.$
In particular, for any fixed $\varepsilon>0$ and any sufficiently large $q\geqslant q_{0}(\varepsilon)$, we have
$F_{q,q-1}(n)\leqslant G_{q,q-1}(n)\leqslant n^{1+\varepsilon}.$
Theorem 1.2 is our main result on this vector problem, asymptotically resolving, as $q\to\infty$, the growth rate of $F_{q,q-1}(n)$ and $G_{q,q-1}(n)$. However, if one is only interested in answering 1.1, there is a much simpler, and quite surprising, argument, which proves the weaker bound
[…]
In fact, (5) follows from the following proposition, which shows that when $r>\frac{2q}{3}$, there is no difference between the $r$-increasing and $r$-comparable questions.
@@
Note that Proposition 1.3 immediately implies (5): for $q\geqslant 4$, we have the chain of inequalities
$G_{q,q-1}(n)=F_{q,q-1}(n)\leqslant F_{3,2}(n)\leqslant n^{2-\delta},$
from Proposition 1.3, (2), and (3), respectively. That is, Proposition 1.3 allows us to reduce to the theorem of Gowers and Long, and gives us an extremely small, but positive, power saving on the trivial upper bound on $G_{q,q-1}(n)$.
Proposition 1.3 is quite surprising, since, as discussed above, it is probably false for $(q,r)=(3,2)$; we have that $G_{3,2}(n)\geqslant n^{1.54}$, whereas Gowers and Long conjectured $F_{3,2}(n)\leqslant n^{3/2}$. This also suggests that the requirement $r>\frac{2q}{3}$ in Proposition 1.3 is best possible. Despite this, the proof of Proposition 1.3 is extremely simple.
@@
1.2. Packing pods in $q$-dimensional space
A tripod of order $n$ consists of a corner and the three adjacent edges of an integer $n\times n\times n$ cube. The corner of the tripod is called its apex. We say that a collection of tripods is aligned if they are translates of one other, and that they pack if they are pairwise disjoint. The maximum number of aligned tripods which pack and whose apex is in $n]^{3}$ is precisely $G_{3,2}(n)$, as two tripods with apices in $n]^{3}$ are disjoint if and only if their apices are $2$-comparable. Recall that Tiskin [37] proved that $G_{3,2}(n)=o(n^{2})$; his motivation came from the question of packing tripods, and he proved the equivalent statement that the density of $3$-space that can be filled with tripods of order of $n$ is $o(1)$ as $n\to\infty$. Indeed, it is not hard to show that the two questions are equivalent: bounding the maximum number of disjoint tripods of order $n$ whose apex lies in $n]^{3}$ is equivalent to bounding the maximum density of $3$-space that can be packed by order-$n$ tripods. For a formal statement (and in particular for the formal definition of the density of a packing), we refer to [35, Chapter 4] or [32].
(a) A tripod of order 4
last coordinate = 1last coordinate = 2last coordinate = 0
As above, it is not hard to convert between the question of packing $(q,r)$-pods of order $n$ with apices in $n]^{q}$ and the question of densely packing $q$-dimensional space with $(q,r)$-pods of order $n$. From Theorem 1.2 we thus get a stronger bound on the density of $q$-space that can be packed by (q,q1)(q,q-1)-pods for larger $q$.
Corollary 1.4.
For $n$ sufficiently large as a function of $q$, the density of $q$-space that can be packed by (q,q1)(q,q-1)-pods of order $n$ is at most $n^{O(1/\sqrt{\log q})-1}$.
Here we used the fact that there are only $2^{q}$ corners of the $q$-cube, so by losing a factor $2^{q}$, we can assume that all the (q,q1)(q,q-1)-pods we pack are aligned.
1.3. Color-avoiding paths
Our proof of Theorem 1.2 uses the language of color-avoiding paths in tournaments. This is both a much more convenient setting in which to place our arguments, and is also a setting with a large number of natural questions. We now turn to discuss these problems and our new results on them.
The kind of question we study, wherein one is given a colored structure, and wishes to find a substructure avoiding one of the colors, has a rich history in Ramsey theory. The study of such questions goes back at least to the seminal work of Erdős and Szemerédi [16], and is a special case of the “generalized Ramsey theory” introduced by Erdős and Gyárfás [14] and intensively studied since. We also refer to the recent paper [28], which studies a color-avoiding question in hypergraphs that is closely related to our main topic.
The most basic question of the kind we study was introduced by Loh [26], who asked for the length of the longest color-avoiding monotone path that is guaranteed in every $3$-coloring of $E(K_{N})$. The trivial lower bound here is $\sqrt{N}$, obtained by simply merging two of the colors and applying the Erdős–Szekeres lemma to find a monochromatic monotone path in the modified coloring; any such path must receive at most two colors in the original coloring, and hence is color-avoiding. On the other hand, the best known upper bound for this problem is $N^{2/3}$, which is obtained by a simple product construction (see Proposition 3.3 for details).
The similarity of these bounds to (1) is not a coincidence. Indeed, Loh observed that one can convert this into a vector problem, by assigning to every vertex $a$ of $K_{N}$ a vector $x_{a}\in\mathbb{N}^{3}$, where the $i$th coordinate of $x_{a}$ is the length of the longest monotone path ending at $a$ which avoids the $i$th color. The key observation is that if $a<b$ and the edge $ab$ receives color $i$, then for both choices of $j\in[3]\setminus\{i\}$, we can extend a longest $j$-avoiding path ending at $a$ to a strictly longer one ending at $b$. This implies that $(x_{a})_{j}<(x_{b})_{j}$ for both choices of $j$, and hence $x_{a}<_{2}x_{b}$. In particular, if the longest color-avoiding monotone path has length at most $n$, then we have found a $2$-increasing sequence of vectors in $n]^{3}$, implying that $N\leqslant F_{2,3}(n)$. As it turns out (see Proposition 3.1), this connection can be reversed, and the study of $F_{3,2}(n)$ is equivalent to the study of the longest color-avoiding monotone path in a $3$-coloring of $E(K_{N})$. More generally, $F_{q,q-1}(n)$ corresponds to finding color-avoiding monotone paths in $q$-colorings of $E(K_{N})$.
As a consequence, one can convert the results discussed previously to this setting. Tiskin’s [37] argument (observed independently by Loh [26] in this context) allows us to improve the trivial bound above, and always find a color-avoiding monotone path of length $\omega(\sqrt{N})$. More strongly, the bound (3) of Gowers and Long [22] implies, for $q>3$, that every $q$-edge-coloring of $E(K_{N})$ contains a color-avoiding path of length $N^{\frac{1}{2}+\delta}$, for some absolute constant $\delta>0$. On the other hand, the best known upper bound for this problem, corresponding to the lower bound on $F_{q,q-1}(n)$, is that there is a $q$-coloring of $E(K_{N})$ whose longest color-avoiding directed path has length $N^{1-1/q}$. The gap between the exponents for large $q$ is equivalent to that in (4). Our main theorem in this setting, which is equivalent to Theorem 1.2, shows that the true exponent tends to $1$ as $q$ grows.
Theorem 1.5.
Let $q\geqslant 4$. In any $q$-coloring of $E(K_{N})$, there exists a color-avoiding monotone path of length at least $N^{1-O(1/\sqrt{\log q})}$.
As we saw in our discussion of the Erdős–Szekeres lemma, statements about monotone paths are equivalent to statements about directed paths in transitive tournaments. More generally, it is thus natural to ask such questions for arbitrary tournaments. Our original motivation to study this more general question came from the study of $r$-comparable sequences, whose structure naturally corresponds to an arbitrary tournament. Proposition 1.3 implies that in the vector setting, there is nothing new in this seemingly more general problem for $r>\frac{2q}{3}$, but in the Ramsey-theoretic setting, working with general tournaments seems to be genuinely harder and more general.
Our second main result is a strengthening of Theorem 1.5, obtaining essentially the same conclusion for all tournaments.
Theorem 1.6.
Let $q\geqslant 4$. In any $q$-edge coloring of any $N$-vertex tournament, there exists a color-avoiding directed path of length at least $c_{q}N^{1-O(1/\sqrt{\log q})}$, where $c_{q}>0$ is a constant depending only on $q$.
Note that the conclusion of Theorem 1.6 is (very) slightly weaker than that of Theorem 1.5, because of the presence of the constant $c_{q}$. When working with transitive tournaments, we can eliminate this constant via a powering trick (very similar to the use of the tensor product trick in the study of Sidorenko’s conjecture, as in e.g. [1, 13]), but this does not appear to be possible when working with arbitrary tournaments.
In the setting of vectors, Proposition 1.3 implies that $F_{q,q-1}(n)=G_{q,q-1}(n)$ for all $q\geqslant 4$, i.e. that the hardest case of the (q1)(q-1)-comparable problem is actually the case when the tournament defined by the vectors is transitive. We do not know how to prove such a result in the setting of color-avoiding paths, nor do we even expect it to be true in general. Nevertheless, we are able to prove an approximate form of Proposition 1.3 in this setting, which says that the problem of color-avoiding directed paths really is much easier if the tournament is far from transitive. To make this precise, let us recall that an $N$-vertex tournament is $\delta$-close to transitive if it can be made transitive by reversing the orientation of at most $\delta N^{2}$ edges, and is $\delta$-far from transitive otherwise.
Theorem 1.7.
If an $N$-vertex tournament is $\delta$-far from transitive, then any $q$-coloring of its edges contains a directed path of length $c\delta^{2}N/q^{3}$ which is colored by at most three colors, where $c>0$ is an absolute constant. Furthermore, the coloring contains the square of a path such that the direction of the edges of consecutive vertices are forward, the direction of edges between vertices of distance two are backwards, and the coloring of the edges is periodic with period $3$.
The proof of Theorem 1.7 is extremely short and simple, and relies on a result of the first two authors [20] on the structure of tournaments that are far from transitive. They proved a directed version of the celebrated triangle removal lemma [30] with very good quantitative bounds, and it is this good quantitative behavior that means that the $\delta$-dependence in Theorem 1.7 is polynomial.
@@ 1.3. Color-avoiding paths
The rest of this paper is organized as follows. We present a high-level outline of the proof of Theorem 1.6 in Section 2. In Section 3, we warm up to the topic by presenting some simple proofs, including the short proof of Theorem 1.7, proofs that the vector and Ramsey notions are equivalent, and some properties of the product construction alluded to previously; in particular, we also show there how Theorem 1.6 implies Theorems 1.2 and 1.5. Finally, we present the full proof in Section 4.
@@ 2. Proof overview
In this section, we give a sketch of the proof of Theorem 1.6. For the majority of our discussion, we focus on the case of transitive tournaments; most of the ideas used in the proof of the full theorem are already present in this setting, with fewer technical complications. At the end of the sketch, we describe which additional ingredients are needed for the general case.
As our tournament is transitive, we can simply view it as an ordered complete graph on the vertex set $[N]$; additionally, a path in this complete graph is directed if and only if it is monotone, i.e. its vertices are strictly increasing as integers in $[N]$. So our task is to prove that any $q$-edge-colored complete graph on $[N]$ contains a color-avoiding monotone path of length at least $N^{1-C/\sqrt{\log q}}$, and we prove this by induction on $N$. For this proof sketch, we will not worry too much about the precise quantitative error term, and only aim to prove a lower bound of the form $N^{1-o(1)}$, for some function $o(1)$ that tends to 0 as $q\to\infty$. For a $q$-edge-colored transitive tournament $T$ and a color $i\in[q]$, let $\ell_{i}(T)$ denote the length of the longest directed path in $T$ avoiding color $i$. Thus, our goal is to prove by induction that $\ell_{i}(T)\geqslant N^{1-o(1)}$ for some $i$. We will actually strengthen the induction hypothesis, and prove that $\Pi(T)\geqslant N^{q-o(q)}$, where we define $\Pi(T)=\prod_{i=1}^{q}\ell_{i}(T)$. The statement $\Pi(T)\geqslant N^{q-o(q)}$ immediately implies that $\ell_{i}(T)\geqslant N^{1-o(1)}$ for some $i\in[q]$, hence this really is a stronger statement.
@@
Concretely, set $s=\varepsilon N/q$, where $\varepsilon>0$ is a small parameter depending only on $q$ (and tending to 0 as $q\to\infty$). We wish to pick $s$ different $i$-avoiding paths in each half; moreover, since our goal is to try gluing every pair of paths from each half, we want these $s$ paths to have different endpoints. To accomplish this, let us define $\ell_{i}(^{\to}v)$ to be the length of the longest $i$-avoiding directed path ending at $v$, and similarly $\ell_{i}(w^{\to})$ to be the length of the longest $i$-avoiding directed path starting at $w$. We now define $X_{i}$ to be the set of the $s$ vertices in the first half with the highest value of $\ell_{i}(^{\to}\bullet)$. Similarly, we define $Y_{i}$ to be the set of the $s$ vertices in the second half with the largest value of $\ell_{i}(\bullet^{\to})$. The paths $P_{1}$ that we consider will then be the longest $i$-avoiding paths ending at a vertex of $X_{i}$, and similarly the $P_{2}$ we consider are the longest $i$-avoiding paths starting at a vertex of $Y_{i}$. Heuristically, since $s\ll N$ and $X_{i}$ consists of the $s$ “best” endpoints for $i$-avoiding paths in $T_{1}$, any path ending at a vertex of $X_{i}$ is the endpoint of such a path of length almost $\ell_{i}(T_{1})$.
@@
Thus, if nearly all colors are compressed, we are “almost” in the dream scenario and we can prove the claim inductively by absorbing all the losses in the $N^{-o(q)}$ factor. It remains to understand what happens when we have a large number, say $2p$, of non-compressed colors; by symmetry, we may assume that at least $p$ of them are left-diffuse, i.e. that the majority of $X_{i}$ for these colors lies in the interval $1,\frac{N}{2}-4s]$. Let $U_{i}\subseteq X_{i}$ be the vertices far from the midpoint, for each of these $p$ left-diffuse colors $i$.
@@
Apart from the computations, the discussion above is an essentially complete sketch of the proof of Theorem 1.6 in the case that $T$ is transitive. We now discuss the difficulties encountered when passing to the setting of general tournaments, and the main additional ideas needed to overcome them.
@@
However, there is a major thing that stops working, which is instrumental to using the key claim above. Namely, if $v$ precedes $w$ in the ordering, and even if $v\to w$ in $T$, we may have that $vw$ does not receive color $i$, and yet not be able to conclude that $\ell_{i}(w)>\ell_{i}(v)$. Indeed, it is possible that all long $i$-avoiding paths ending at $v$ have already passed through $w$, and hence we cannot simply extend such a path to end at $w$. This is a fundamental issue [^2] We remark that this issue is quite similar to the main difficulty in proving the Gallai–Hasse–Roy–Vitaver theorem. which is not present in the fully transitive setting: there, a directed path must be monotone, and hence cannot have already used a “future” vertex.
In order to overcome this issue, we split the set of colors into long colors—those colors $i$ where there is an $i$-avoiding path of length at least $\gamma N$, for some small $\gamma>0$—and short colors. For short colors, the arguments used above can now be made to work: the longest $i$-avoiding path ending at $w$ has length at most $\gamma N$, hence there are relatively few vertices $w$ for which we cannot argue about the color of the edge $vw$, as most vertices $w$ are not on such a path. Dealing with these “exceptional” vertices now requires some care, but no genuinely new ideas. On the other hand, we know essentially nothing about the long colors, except that they are long. If we only aimed to prove that $T$ contains a long color-avoiding directed path, then long colors would immediately help us; unfortunately, we strengthened the inductive hypothesis to argue about $\Pi(T)$, so we cannot simply say we are done if there is a long color.

### v2

@@ 1.2. Applications: vectors sequences and pod packings
Thanks to (3), our lower bound on $g_{q,q-1}(N)$ in Theorem 1.1 yields a corresponding upper bound on $G_{q,q-1}(n)$, and thus also on the packing density of (q,q1)(q,q-1)-pods. As such, we are able to resolve 1.3, as stated in the following corollary.
Corollary 1.4.
We have
$G_{q,q-1}(n)\leqslant C_{q}n^{1+O(1/\sqrt{\log q})},$
where $C_{q}>0$ is an absolute constant.
Consequently, for $n$ sufficiently large as a function of $q$, the density of $q$-space that can be packed by (q,q1)(q,q-1)-pods of order $n$ is at most $n^{O(1/\sqrt{\log q})-1}$.
In the second statement, we used the fact that there are only $2^{q}$ corners of the $q$-cube, so by losing a factor $2^{q}$, we can assume that all the (q,q1)(q,q-1)-pods we pack are aligned.
However, if one is only interested in these geometric corollaries, there is an alternative approach that yields much stronger bounds. The first step of the approach is the following surprisingly simple observation, which shows that when $r>\frac{2q}{3}$, there is no difference between the $r$-increasing and $r$-comparable questions.
@@
Proposition 1.5 is quite surprising for a few reasons. First, it is false for $(q,r)=(3,2)$: the old construction of Hamaker and Stein shows $G_{3,2}(7)\geqslant 19$, whereas one can check by computer that $F_{3,2}(7)=17$. More generally, Gowers and Long proved that $G_{3,2}(n)\geqslant n^{1.54}$, and they conjectured that $F_{3,2}(n)\leqslant n^{3/2}$. These facts also suggest that the requirement $r>\frac{2q}{3}$ in Proposition 1.5 is best possible.
The other reason why Proposition 1.5 is surprising to us is that we do not expect it to be true for tournaments, i.e. we do not expect $f_{q,q-1}(N)=g_{q,q-1}(N)$; although Theorem 1.2 shows that the extremal examples for $g_{q,q-1}$ must be close to transitive, we do not expect them to be exactly transitive. Despite all of this, the proof of Proposition 1.5 is extremely simple.
@@
Proposition 1.5 is already enough to answer 1.3, when combined with the result $F_{q,q-1}(n)\leqslant n^{2-\delta}$ of Gowers and Long. However, as pointed out to us by Cosmin Pohoata (personal communication), much stronger results were obtained by Sawin and Tao [35], who proved the following theorem.
Proposition 1.6 (Sawin–Tao).
Let $\gamma\in(\frac{1}{2},1)$, and let $(X_{1},\dots,X_{N})$ be a sequence of (dependent) random variables valued in $[n]$ with the property that $\mathbb{P}(X_{b}>X_{a})\geqslant\gamma$ for all $1\leqslant a<b\leqslant N$. Then
$N\leqslant n^{1/(2\gamma-1)}.$
As noted by Pohoata, this immediately implies that for all $q\geqslant r\geqslant 1$ with $r>\frac{q}{2}$, we have
$F_{q,r}(n)\leqslant n^{\frac{q}{2r-q}}.$ (4)
Indeed, suppose that $x_{1},\dots,x_{N}\in[n]^{q}$ is an $r$-increasing sequence of vectors with $N=F_{q,r}(n)$. Let $Y$ be a uniformly random element of $[q]$, and let $X_{a}$ equal the $Y$th coordinate of $x_{a}$, for all $a$. Then $(X_{1},\dots,X_{n})$ is a sequence of (dependent) random variables valued in $[n]$, and the $r$-increasing condition implies that $\mathbb{P}(X_{b}>X_{a})\geqslant r/q$ for all $a<b$, since $x_{b}$ is strictly larger than $x_{a}$ in at least $r$ of its $q$ coordinates. We may thus apply Proposition 1.6 to conclude that $F_{q,r}(n)=N\leqslant n^{1/(2(r/q)-1)}$, yielding (4). In particular, applying this bound with r=q1r=q-1, we find that
$F_{q,q-1}(n)\leqslant n^{1+\frac{2}{q-2}}$
for all $n$. When $q\geqslant 4$, we can combine this with Proposition 1.5 to conclude that $G_{q,q-1}(n)\leqslant n^{1+\frac{2}{q-2}}$ as well, yielding a much stronger dependence on $q$ than given in Corollary 1.4.
The proof of Proposition 1.6 uses Fourier-analytic methods, and as such seems restricted to the setting of $r$-increasing vector sequences. Thanks to the equivalence (2), we obtain from it a strong bound on $f_{q,q-1}(N)$ as well. However, we do not see how to extend these analytic techniques to handle more general tournaments, and in particular we do not see how to deduce any bound on $g_{q,q-1}(N)$ using these techniques. Thus, while Theorem 1.1 yields weaker bounds in these applications, it appears to be a genuinely more general result.
@@
The rest of this paper is organized as follows. We present a high-level outline of the proof of Theorem 1.1 in Section 2. In Section 3, we warm up to the topic by presenting some simple proofs, including the short proof of Theorem 1.2 and some properties of the product construction alluded to previously. We present the proof of Theorem 1.1 in Section 4, and end in Section 5 with some concluding remarks.
@@ 2. Proof overview
In this section, we give a sketch of the proof of Theorem 1.1. For the majority of our discussion, we focus on the case of transitive tournaments; most of the ideas used in the proof of the full theorem are already present in this setting, with fewer technical complications. Although the Fourier-analytic techniques of Sawin and Tao yield stronger bounds in the transitive setting, they do not seem to extend to the general case, in contrast to the combinatorial techniques we use. At the end of the sketch, we describe which additional ingredients are needed for the general case.
As our tournament is transitive, we can simply view it as an ordered complete graph on the vertex set $[N]$; additionally, a path in this complete graph is directed if and only if it is monotone, i.e. its vertices are strictly increasing as integers in $[N]$. So our task is to prove that any $q$-edge-colored complete graph on $[N]$ contains a color-avoiding monotone path of length at least $c_{q}N^{1-C/\sqrt{\log q}}$, and we prove this by induction on $N$. For this proof sketch, we will not worry too much about the precise quantitative error term, and only aim to prove a lower bound of the form $N^{1-o(1)}$, for some function $o(1)$ that tends to 0 as $q\to\infty$. For a $q$-edge-colored transitive tournament $T$ and a color $i\in[q]$, let $\ell_{i}(T)$ denote the length of the longest directed path in $T$ avoiding color $i$. Thus, our goal is to prove by induction that $\ell_{i}(T)\geqslant N^{1-o(1)}$ for some $i$. We will actually strengthen the induction hypothesis, and prove that $\Pi(T)\geqslant N^{q-o(q)}$, where we define $\Pi(T)=\prod_{i=1}^{q}\ell_{i}(T)$. The statement $\Pi(T)\geqslant N^{q-o(q)}$ immediately implies that $\ell_{i}(T)\geqslant N^{1-o(1)}$ for some $i\in[q]$, hence this really is a stronger statement.
@@
Concretely, set $s=\varepsilon N/q$, where $\varepsilon>0$ is a small parameter depending only on $q$ (and tending to 0 as $q\to\infty$). We wish to pick $s$ different $i$-avoiding paths in each half; moreover, since our goal is to try gluing every pair of paths from each half, we want these $s$ paths to have different endpoints. To accomplish this, let us define $\ell_{i}(^{\to}v)$ to be the length of the longest $i$-avoiding directed path ending at $v$, and similarly $\ell_{i}(w^{\to})$ to be the length of the longest $i$-avoiding directed path starting at $w$. We now define $X_{i}$ to be the set of the $s$ vertices in the first half with the highest value of $\ell_{i}(^{\to}\bullet)$. Similarly, we define $Y_{i}$ to be the set of the $s$ vertices in the second half with the highest value of $\ell_{i}(\bullet^{\to})$. The paths $P_{1}$ that we consider will then be the longest $i$-avoiding paths ending at a vertex of $X_{i}$, and similarly the $P_{2}$ we consider are the longest $i$-avoiding paths starting at a vertex of $Y_{i}$. Heuristically, since $s\ll N$ and $X_{i}$ consists of the $s$ “best” endpoints for $i$-avoiding paths in $T_{1}$, every vertex of $X_{i}$ is the endpoint of such a path of length almost $\ell_{i}(T_{1})$.
@@
Thus, if nearly all colors are compressed, we are “almost” in the dream scenario and we can prove the claim inductively by absorbing all the losses in the $N^{-o(q)}$ factor. It remains to understand what happens when we have a large number, say $2p$, of non-compressed colors; by symmetry, we may assume that at least $p$ of them are left-diffuse, i.e. that the majority of $X_{i}$ for these colors lies in the interval $1,\frac{N}{2}-4s]$. Let $U_{i}=X_{i}\cap[1,\frac{N}{2}-4s]$, for each of these $p$ left-diffuse colors $i$.
@@
Apart from the computations, the discussion above is an essentially complete sketch of the proof of Theorem 1.1 in the case that $T$ is transitive. We now discuss the difficulties encountered when passing to the general setting, and the main additional ideas needed to overcome them.
@@
However, there is a major thing that stops working, which is instrumental to using the key claim above. Namely, if $v$ precedes $w$ in the ordering, and even if $v\to w$ in $T$, we may have that $vw$ does not receive color $i$, and yet not be able to conclude that $\ell_{i}(w)>\ell_{i}(v)$. Indeed, it is possible that all long $i$-avoiding paths ending at $v$ have already passed through $w$, and hence we cannot simply extend such a path to end at $w$. This is a fundamental issue which is not present in the fully transitive setting: there, a directed path must be monotone, and hence cannot have already used a “future” vertex.
In order to overcome this issue, we split the set of colors into long colors—those colors $i$ where there is an $i$-avoiding path of length at least $\gamma N$, for some small $\gamma>0$—and short colors. For short colors, the arguments used above can now be made to work: the longest $i$-avoiding path ending at $v$ has length at most $\gamma N$, hence there are relatively few vertices $w$ for which we cannot argue about the color of the edge $vw$, as most vertices $w$ are not on such a path. Dealing with these “exceptional” vertices now requires some care, but no genuinely new ideas. On the other hand, we know essentially nothing about the long colors, except that they are long. If we only aimed to prove that $T$ contains a long color-avoiding directed path, then long colors would immediately help us; unfortunately, we strengthened the inductive hypothesis to argue about $\Pi(T)$, so we cannot simply say we are done if there is a long color.